Find T(x,t) in a laterally insulated 2 m-long rod if k=10-4 m2/s and T(x,0)=100(2x-x2), T(0,t)=0=T(2,t).
1
Expert's answer
2019-11-01T10:14:28-0400

"T(0,t)=0=T(2,t)" ----- (i)

This is the boundary condition for laterally insulated rod having ends maintained at certain temperatures.

Governing equation for the following condition

"\u2202T(x,t)\/\u2202t=k \u2202^2T(x,t)\/\u2202x^2" ------------- (ii)

For separation of variable "T(x,t) = f(x).g(t)" ---------- (iii)

Substituting equation (iii) in equation (ii);

we get "g'(t)\/g(t) = k f''(x)\/f(x) =l" ----- (iv)

"g'(t)\/g(t)=l"

"d(g(t))\/g(t)=ldt"

Integrating both sides, we get

"ln (g(t)\/c_2) = lt"

"g(t)\/c_2 = exp(lt)"

"g(t) = c_2 exp(lt)" ----(v)

"k f''(x)\/f(x) =l"

"\u2202^2f(x)\/\u2202x^2=(l\/k)f(x)"

"\u2202^2f(x)\/\u2202x^2 =-m^2f(x)"

"l\/k= -m^2" -----(vi)

Solution to this differential equation is

"f(x) = c_1sin(mx)+c_2 cos(mx)"

"T(x,t) =(c_2 exp(lt))(c_1sin(mx)+c_3cos(mx))"


Using equation (i)

"T(0,t) =(c_2 exp(lt))(c_1sin(0)+c_3cos(0))= 0"

"T(0,t) =c_3c_2 exp(lt)= 0"

"c_3 =0"

Using equation (i)

"T(2,t)=(c_2 exp(lt))(c_1sin(m.2)) =0"

"T(2,t)=c_1c_2exp(lt)sin(m.2)=0"

We obtain "2.m= n\u03c0"

"m=n\u03c0\/2"

Using equation (vi)

"l\/k=-m^2=-n^2\u03c0^2\/4"

"l =-n^2\u03c0^2k\/4"

So the equation becomes

"T(x,t) =\sum_{n=1}^{\infty}(c_2 exp(-n^2π^2kt/4))(c_1sin(nπx/2))"

"T(x,t) =\sum_{n=1}^{\infty}(c_n(sin(nπx/2) exp(-n^2π^2kt/4))"

"c_n= c_1.c_2"

We have to find cn now

"T(x,0) =\sum_{n=1}^{\infty}(c_n(sin(nπx/2) exp(-n^2π^2k.0/4))"

"\sum_{n=1}^{\infty} (c_n(sin(nπx/2) = 100(2x-x^2)"

It is a sine Fourier series.we know;

"c_n = 2×\int_{0}^{2}(100(2x-x^2)(sin(nπx/2))dx"

solving the integration using by parts and then putting the limits we get,

"c_n = -3200\/(n^3\u03c0^3)((-1)^n-1)"

Answer:

"T(x,t) = \sum_{n=1}^{\infty} (c_n(sin(nπx/2) exp(-n^2π^2kt/4))",

where "c_n = -3200\/(n^3\u03c0^3)((-1)^n-1)"